OpenStudy (anonymous):

find the term that contains x^8 y^9 in the expansion of (x^2+3y^3)^7

Parth (parthkohli):

first term in the expansion:\[\binom{7}{0}\left(x^{2}\right)^7\left(3y^3\right)^0 = x^{14}\] second:\[\binom{7}{1}\left(x^2\right)^6 \left(3y^3\right)^1 = 21x^{14}y^3\]Seems that the one which brings the 9th power is where \(3y^3\) is raised to the third. So the fourth term. Find out what that is.

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

help me tooo apex stuff